If 10.0 g of a solute is added to 100.0 g of water, the mass percent of the solution is



1
Expert's answer
2021-11-15T11:00:07-0500

Mass of solute "+" water


"=100+10"

"=110g"


"\boxed {\textcolor {blue}{\%m/m=\frac{gram \>of\> solute}{mass\> of\> solution}×100\%}}"


"\frac{10}{110}×100\%=9.091\%"